# XXLTraffic: Expanding and Extremely Long Traffic Dataset for Ultra-Dynamic Forecasting Challenges

### v1

@@ 4.3 Data Overview
Figure 3: XXLTraffic dataset overview and its evolving development. This figure provides a global overview and two local overviews, showcasing the diversity of sensor distribution. The lower parts highlights a selected region to illustrate the growth and changes in traffic sensors over time.
It is evident that sensors at the same location may collect completely different distributions over the course of urban evolution. As shown in Figure 4, some sensors have maintained the same distribution from 2005 to 2024, while others have experienced significant changes in distribution. The temporal changes causing domain shifts present a significant challenge for our extremely long forecasting.
Figure 4: Sensor traffic status distribution of District 8 from 2005 to 2024. While some sensors exhibit minimal changes, others show significant distribution differences, regardless of whether they are in low-traffic or high-traffic areas. This presents substantial challenges for extremely long forecasting with long gaps.
4.4 XXLTraffic Licence
@@ 5 Experiments
We conducted experiments for both extremely long forecasting with gaps using gap dataset and conventional long-term forecasting using hourly dataset and daily dataset. Additionally, referring to the definition in Figure 2, we set the gap parameter $g$ as 1 year, 1.5 years, and 2 years for the gap dataset, as illustrated by Figure 5.
@@
Figure 5: Problem definition. The yellow boxes represent typical predictions, the gray boxes denote gap periods between observation and prediction, and the blue boxes indicate extended predictions.
5.1 Datasets
@@
We conducted experiments on all proposed sub-datasets. To maintain consistency with previous state-of-the-art benchmarks, we selected districts 03, 04, 07, and 08 (widely recognized as PEMS03/04/07/08) for our experiments, using the gap, hourly, and daily datasets from these districts. Results for other datasets are presented in Appendix. All sub-datasets were divided into training, validation, and test sets using a 6:2:2 ratio. For the gap dataset, due to the extensive span of up to 20 years resulting in a large sample size, we fixed a seed during data preprocessing to select 10% of the dataset for training and testing to quickly demonstrate our results. In our experiments, due to GPU memory constraints, we pruned PEMS07_ours and PEMS12_ours by removing nodes with a missing data ratio greater than 10%, reducing the node number to 1,613 and 867. The details of the datasets used in our benchmarking is in Appendix A.1.
Table 3: Comparison in gap dataset. The bold text indicates the best.
Methods Mamba iTransformer DLinear Autoformer
Metrics MSE MAE MSE MAE MSE MAE MSE MAE
PEMS03_gap 1-year gap 96 1.457 0.913 1.636 0.989 1.500 0.933 1.301 0.906
192 1.472 0.922 1.597 0.970 1.542 0.945 1.266 0.867
336 1.434 0.913 1.512 0.935 1.531 0.935 1.137 0.807
1.5-year gap 96 1.485 0.942 1.879 1.078 1.653 0.976 1.467 0.945
192 1.442 0.934 1.753 1.017 1.642 0.968 1.460 0.923
336 1.446 0.938 1.662 0.987 1.632 0.968 1.345 0.882
2-year gap 96 1.359 0.833 1.844 1.048 1.568 0.954 1.328 0.894
192 1.216 0.772 1.729 1.008 1.473 0.896 1.235 0.837
336 1.294 0.801 1.614 0.960 1.407 0.870 1.220 0.838
PEMS04_gap 1-year gap 96 1.325 0.914 1.644 1.037 1.396 0.955 0.819 0.717
192 1.438 0.642 1.587 1.018 1.440 0.965 0.941 0.767
336 1.424 0.935 1.447 0.960 1.421 0.954 0.853 0.730
1.5-year gap 96 1.204 0.890 2.014 1.193 1.488 0.995 0.981 0.779
192 0.982 0.787 1.649 1.038 1.301 0.913 0.867 0.735
336 0.961 0.792 1.352 0.929 1.298 0.908 0.762 0.679
2-year gap 96 1.220 0.893 1.652 1.074 1.446 0.977 0.909 0.755
192 1.004 0.807 1.189 0.870 1.268 0.893 0.909 0.747
336 1.198 0.864 1.584 1.016 1.269 0.898 0.898 0.736
192 1.637 0.996 1.816 1.061 1.776 1.043 1.202 0.852
336 1.637 0.991 1.784 1.028 1.774 1.034 1.098 0.774
1.5-year gap 96 1.571 0.972 1.705 1.016 1.585 0.987 1.184 0.844
192 1.578 0.982 1.657 1.004 1.624 0.996 1.088 0.800
336 1.818 1.127 1.712 1.028 1.668 1.036 0.991 0.729
2-year gap 96 5.411 1.319 2.055 1.149 1.746 1.048 1.099 0.802
192 12.620 1.370 2.006 1.128 1.705 1.002 1.284 0.875
336 9.614 1.378 1.812 1.057 1.605 0.968 0.972 0.734
PEMS08_gap 1-year gap 96 5.411 1.319 1.514 0.950 1.771 1.058 1.153 0.843
192 12.620 1.370 1.499 0.900 1.762 1.059 1.202 0.845
336 9.614 1.378 1.654 0.984 1.961 1.131 1.184 0.849
1.5-year gap 96 4.453 1.061 2.286 1.196 1.978 1.072 1.428 0.901
192 9.413 1.046 1.713 0.971 1.606 0.928 1.314 0.833
336 10.457 1.063 1.890 1.488 1.736 0.985 1.320 0.836
2-year gap 96 5.117 0.953 2.400 1.232 1.969 1.069 1.474 0.895
192 12.769 0.933 1.974 1.070 1.670 0.942 1.505 0.883
336 13.382 0.946 1.936 1.056 1.628 0.926 1.464 0.874
5.2 Baselines
@@
In our comparison experiments, we adopted four popular baselines, including MLP, Transformer, and Mamba architectures. Autoformer [37], an earlier state-of-the-art model, leverages a decomposition architecture and auto-correlation mechanism to enhance efficiency and accuracy in long-term time series forecasting, outperforming traditional Transformer models. iTransformer [26] is the latest and most effective Transformer-based model, utilizing attention and feed-forward networks on inverted dimensions, embedding time points into variate tokens. DLinear [43] challenges the effectiveness of Transformer models by proposing a simple one-layer linear model that captures temporal relations in an ordered set of continuous points. It employs positional encoding and uses tokens to embed sub-series, preserving some ordering information in Transformers. Lastly, Mamba [8], a well-known sequential model from last year, uses a bidirectional Mamba block to extract inter-variate correlations and temporal dependencies.
5.3 Implementation Details
@@
We adopted the default settings of the Time-Series-Library [36] to conduct a comprehensive comparison of baselines. We use the results of five random seeds as the average. We used 96 time steps as input and 336 time steps as ground truth. The code was implemented in PyTorch and executed on a V100 GPU with 32GB memory and 384GB RAM, provided by NCI Australia, an NCRIS-enabled capability supported by the Australian Government.
5.4 Results of Extremely Long Forecasting with gaps
@@
We use Mean Squared Error (MSE) and Mean Absolute Error (MAE) metrics to evaluate performance, averaging results across different seeds. It is observed that nearly all results are poor, highlighting the significant challenge posed by domain shifts over time for extremely long forecasting with gaps. These baseline results also indicate that traditional SOTA rankings and methodologies are no longer effective. Notably, Autoformer, which performs the worst under conventional data and settings, shows the best performance in our setting. This is understandable because Autoformer’s core technology focuses on exploring the correlation within the data itself. This insight suggests that future efforts to tackle this problem should place greater emphasis on leveraging the intrinsic potential of the data.
@@
Table 4: Results of ablation study with 4 different input step lengths
@@
5.5 Results of Hourly and Daily Forecasting
We believe that both the hourly and daily datasets are equally significant. Multi-scale, diverse datasets can provide the community with valuable references. We observe that the performance degrades progressively from the hourly to the daily to the gap datasets. Smaller time scales help reduce complexity and uncertainty, thereby improving prediction accuracy. Research has shown that clustering at different scales can enhance model performance [34]. Therefore, our aggregated version of the data will contribute new external features to the community.
Table 5: Comparison in hourly and daily datasets
@@ 5.5 Results of Hourly and Daily Forecasting
6 Prospects and Constraints
Prospects. Our dataset spans the longest time period among existing datasets, and it is not only the largest spatially, but also evolving in growth. It can continue to update in line with the updates from the PeMS system in the future. It is specifically designed for various complex scenarios, such as those already mentioned, including extremely long forecasting with long gaps, and hourly and daily predictions. Additionally, zero-shot forecasting designed for evolving growth scenarios will also be included in Appendix.
@@ 6 Prospects and Constraints
Acknowledgments and Disclosure of Funding
We would like to acknowledge the support of Cisco’s National Industry Innovation Network (NIIN) Research Chair Program and the ARC Centre of Excellence for Automated Decision-Making and Society (CE200100005). We acknowledge the resources and services from the National Computational Infrastructure (NCI), which is supported by the Australian Government.
Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes] See Section 1
    2. (b)
      Did you describe the limitations of your work? [Yes] See Section 6
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [Yes] We have discussed and identified no negative societal impact.
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] See Section 4.1
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [Yes] See Section 5.1 and Section 5.3
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes] As discussed in Section 5.3, our experiments area averaged over five random seeds.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] See Section 5.3
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [Yes] We have cited the creators in Section 4.1
    2. (b)
      Did you mention the license of the assets? [Yes] See Section 4.4. Our dataset complies with the CC BY-NC 4.0 license, and the code adheres to the MIT license.
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [No]
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] Please refer to Section 4.1 for the detailed discussion.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [No] This dataset is distributed across the highway, collecting traffic speed, traffic flow, and traffic occupancy. It does not contain personally identifiable information or offensive content.
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [N/A]
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [N/A]

### v2

@@ 4.2. Data Preprocessing
Figure 4. XXLTraffic dataset overview and its evolving development. This figure provides a global overview and two local overviews, showcasing the diversity of sensor distribution. The lower parts highlights a selected region to illustrate the growth and changes in traffic sensors over time.
4.3. Data Overview
The XXLTraffic dataset is distributed across highways in the state of California, as illustrated in the Fig. 4(a). The nine colors represent nine districts. From Figures. 4(b),  4(c), and 4(d), we can clearly observe the evolutionary growth of the sensors. The sensors are extensively distributed across both urban and suburban areas, offering diverse modalities. Additionally, the sensors are densely interconnected, enabling the formation of a high-quality traffic graph dataset.
It is evident that sensors at the same location may collect completely different distributions over the course of urban evolution. As shown in Fig. 5, some sensors have maintained the same distribution from 2005 to 2024, while others have experienced significant changes in distribution. The temporal changes causing domain shifts present a significant challenge for our extremely long forecasting.
(a) distribution of District 8 in PeMS.
(b) distribution of tfNSW.
Figure 5. Sensor traffic status distribution of District 8 in PeMS from 2005 to 2024 in 5(a) and from 2016 to 2022 in NSW in 5(b). While some sensors exhibit minimal changes, others show significant distribution differences, regardless of whether they are in low-traffic or high-traffic areas. This presents substantial challenges for extremely long forecasting with long gaps.
4.4. XXLTraffic Licence
@@ 5. Experiments
We conducted experiments for both extremely long forecasting with gaps using gap dataset and conventional long-term forecasting using hourly dataset and daily dataset. Additionally, referring to the definition in Fig. 3, we set the gap parameter $g$ as 1 year, 1.5 years, and 2 years for the gap dataset, as illustrated by Fig. 6.
@@
Figure 6. Problem definition. The yellow boxes represent typical predictions, the gray boxes denote gap periods between observation and prediction, and the blue boxes indicate extended predictions.
5.1. Datasets
@@
We conducted experiments on all proposed sub-datasets. To maintain consistency with previous state-of-the-art benchmarks, we selected districts 03, 04, and 08 (widely recognized as PEMS03/04/08) for the experiments using the gap, and districts 03, 04,07 and 08 for hourly, and daily experiments. Results for other datasets are presented in Appendix. All sub-datasets were divided into training, validation, and test sets using a 6:2:2 ratio. For the gap dataset, due to the extensive span of up to 20 years resulting in a large sample size, we fixed a seed during data preprocessing to select 10% of the dataset for training and testing to quickly demonstrate our results. The details of the datasets used in our benchmarking is in Appendix A.1.
Table 3. Comparison in gap dataset. The bold text indicates the best.
Gap Data Gap Metric Horizon Mamba iTrans DLinear Autofor Infor FEDFor MICN PatchTST STGCN ASTGCN GWN AGCRN PDFor
PEMS03 1-year MSE 96 1.457 1.636 1.500 1.301 0.673 0.934 0.514 0.803 0.556 0.765 0.676 0.596 0.621
192 1.472 1.597 1.542 1.266 0.739 0.960 0.528 0.852 0.562 0.764 0.581 0.565 0.545
336 1.434 1.512 1.531 1.137 0.699 0.849 0.493 0.825 0.561 0.717 0.582 0.580 0.574
MAE 96 0.913 0.989 0.933 0.906 0.552 0.697 0.515 0.647 0.536 0.618 0.574 0.562 0.576
192 0.922 0.970 0.945 0.867 0.581 0.715 0.513 0.664 0.539 0.626 0.546 0.543 0.536
336 0.913 0.935 0.935 0.807 0.560 0.640 0.491 0.636 0.538 0.598 0.543 0.552 0.548
1.5-year MSE 96 1.485 1.879 1.653 1.467 1.138 1.190 0.839 1.245 1.256 1.441 1.250 1.168 1.256
192 1.442 1.753 1.642 1.266 1.276 1.374 0.844 1.339 1.187 1.395 1.194 1.122 1.117
336 1.446 1.662 1.632 1.137 1.340 1.259 0.755 1.359 1.167 1.273 1.015 1.176 1.182
MAE 96 0.942 1.078 0.976 0.945 0.763 0.795 0.686 0.863 0.884 0.959 0.875 0.866 0.899
192 0.934 1.017 0.968 0.867 0.827 0.715 0.681 0.906 0.872 0.947 0.856 0.843 0.842
336 0.938 0.987 0.968 0.807 0.845 0.640 0.632 0.903 0.853 0.877 0.832 0.862 0.871
2-year MSE 96 1.359 1.844 1.568 1.328 1.642 1.205 1.359 1.817 2.059 2.236 1.987 1.950 2.068
192 1.216 1.729 1.473 1.235 1.955 1.489 1.308 1.899 2.055 2.180 1.877 1.746 1.897
336 1.294 1.614 1.407 1.220 1.982 1.615 0.966 1.656 2.022 1.889 1.847 1.930 1.982
MAE 96 0.833 1.048 0.954 0.894 0.956 0.816 0.911 1.104 1.176 1.234 1.148 1.150 1.189
192 0.772 1.008 0.896 0.837 1.067 0.933 0.866 1.119 1.184 1.213 1.117 1.070 1.131
336 0.801 0.960 0.870 0.838 1.086 0.978 0.700 1.016 1.169 1.117 1.107 1.140 1.161
PEMS04 1-year MSE 96 1.325 1.644 1.396 0.819 0.624 0.712 0.721 1.309 1.233 1.311 1.310 1.290 1.101
192 1.438 1.587 1.440 0.941 0.694 0.679 0.611 1.358 1.398 1.443 1.255 1.318 1.078
336 1.424 1.447 1.421 0.853 0.668 0.584 0.526 1.338 1.505 1.325 1.292 1.302 1.085
MAE 96 0.914 1.037 0.955 0.717 0.593 0.650 0.632 0.917 0.904 0.929 0.870 0.865 0.851
192 0.642 1.018 0.965 0.767 0.634 0.638 0.586 0.932 1.025 1.023 0.879 0.907 0.816
336 0.935 0.960 0.954 0.730 0.615 0.592 0.534 0.916 1.104 0.938 0.865 0.899 0.821
1.5-year MSE 96 1.204 2.014 1.488 0.981 0.618 0.664 0.638 1.171 1.369 1.350 1.684 1.222 1.097
192 0.982 1.649 1.301 0.867 0.622 0.679 0.570 1.109 1.642 1.549 1.501 1.346 1.214
336 0.961 1.352 1.298 0.762 0.646 0.488 0.482 1.158 1.368 1.199 1.584 0.231 1.130
MAE 96 0.890 1.193 0.995 0.779 0.592 0.632 0.581 0.870 1.082 0.937 1.063 0.961 0.849
192 0.787 1.038 0.913 0.735 0.588 0.638 0.547 0.846 1.035 1.075 0.988 0.989 0.903
336 0.792 0.929 0.908 0.679 0.598 0.527 0.494 0.850 0.862 0.832 0.997 0.963 0.854
2-year MSE 96 1.220 1.652 1.446 0.909 0.650 0.685 0.666 1.284 1.653 1.247 1.669 1.236 1.099
192 1.004 1.189 1.268 0.909 0.639 0.621 0.596 1.151 1.545 1.356 1.554 1.269 1.159
336 1.198 1.584 1.269 0.898 0.717 0.521 0.481 1.205 1.556 1.174 1.128 1.314 1.032
MAE 96 0.893 1.074 0.977 0.755 0.604 0.644 0.609 0.913 1.074 0.894 1.057 0.901 0.861
192 0.807 0.870 0.893 0.747 0.601 0.607 0.570 0.867 1.023 0.948 1.041 0.915 0.891
336 0.864 1.016 0.898 0.736 0.640 0.552 0.502 0.881 1.011 0.842 0.869 0.947 0.816
PEMS08 1-year MSE 96 5.411 1.514 1.771 1.153 - 1.636 1.556 1.926 2.531 3.119 2.185 2.158 1.921
192 12.620 1.499 1.762 1.202 - 1.401 1.408 1.887 2.560 2.405 2.223 2.144 2.248
336 9.614 1.654 1.961 1.184 - 1.870 1.093 1.962 2.482 2.086 2.228 2.036 1.994
MAE 96 1.319 0.950 1.058 0.843 - 1.384 0.979 1.104 1.325 1.404 1.280 1.220 1.148
192 1.370 0.900 1.059 0.845 - 0.877 0.918 1.106 1.337 1.276 1.311 1.213 1.260
336 1.378 0.984 1.131 0.849 - 1.078 0.761 1.133 1.317 1.181 1.312 1.185 1.179
1.5-year MSE 96 4.453 2.286 1.978 1.428 - 1.389 1.034 1.362 1.772 2.370 1.309 1.166 1.664
192 9.413 1.713 1.606 1.314 - 1.179 1.039 1.666 1.522 1.730 1.074 1.144 1.518
336 10.457 1.890 1.736 1.320 - 1.197 0.868 1.272 1.495 1.444 1.533 1.078 1.049
MAE 96 1.061 1.196 1.072 0.901 - 0.871 0.758 0.867 1.084 1.192 0.896 0.808 1.055
192 1.046 0.971 0.928 0.833 - 0.768 0.755 1.001 0.979 1.043 0.805 0.810 1.012
336 1.063 1.488 0.985 0.836 - 0.782 0.680 0.853 0.966 0.948 0.999 0.784 0.801
2-year MSE 96 5.117 2.400 1.969 1.474 - 1.494 1.030 1.393 1.336 1.962 1.789 1.393 1.071
192 12.769 1.974 1.670 1.505 - 1.292 1.044 1.229 1.218 1.292 1.007 1.229 1.171
336 13.382 1.936 1.628 1.464 - 1.253 0.904 1.246 1.181 1.227 1.181 1.246 1.003
MAE 96 0.953 1.232 1.069 0.895 - 0.885 0.740 0.856 0.868 1.069 1.017 0.856 0.792
192 0.933 1.070 0.942 0.883 - 0.795 0.741 0.813 0.819 0.795 0.761 0.813 0.826
336 0.946 1.056 0.926 0.874 - 0.772 0.685 0.814 0.806 0.842 0.813 0.814 0.768
tfNSW 1-year MSE 96 1.480 1.236 1.166 1.320 1.354 1.377 1.219 0.946 1.176 1.404 1.451 1.283 1.070
192 1.543 1.088 1.184 1.599 1.229 1.312 1.262 1.026 1.495 1.547 1.580 1.267 1.019
336 1.459 1.099 1.196 1.477 1.242 1.219 1.366 0.949 1.517 1.617 1.534 1.048 0.822
MAE 96 0.845 0.822 0.748 0.879 0.777 0.895 0.778 0.778 0.776 0.874 0.871 0.824 0.756
192 0.867 0.733 0.758 0.961 0.757 0.890 0.799 0.821 0.880 0.889 0.904 0.828 0.753
336 0.847 0.739 0.763 0.934 0.776 0.838 0.833 0.774 0.901 0.910 0.892 0.786 0.710
1.5-year MSE 96 1.745 1.355 1.372 1.658 1.615 1.440 1.204 1.025 1.220 1.286 1.598 1.249 1.153
192 1.784 1.260 1.367 1.666 1.484 1.452 1.251 1.058 1.624 1.615 1.620 1.228 1.095
336 1.628 1.284 1.385 1.515 1.500 1.346 1.203 0.995 1.621 1.621 1.220 1.163 1.049
MAE 96 0.944 0.877 0.846 1.016 0.879 0.919 0.778 0.807 0.785 0.813 0.940 0.809 0.790
192 0.968 0.818 0.846 0.997 0.854 0.932 0.789 0.823 0.934 0.902 0.950 0.808 0.792
336 0.939 0.835 0.854 0.945 0.883 0.871 0.770 0.788 0.945 0.900 0.814 0.811 0.775
2-year MSE 96 1.364 1.019 1.008 1.319 1.101 1.318 1.033 0.848 1.201 1.331 1.414 1.124 0.969
192 1.195 0.884 1.024 1.319 1.054 1.222 1.094 0.909 1.287 1.061 1.109 1.058 0.941
336 1.588 0.907 1.166 1.233 0.787 1.056 0.959 0.865 1.260 0.996 1.011 0.960 0.917
MAE 96 0.857 0.782 0.742 0.907 0.713 0.879 0.745 0.727 0.785 0.823 0.862 0.749 0.702
192 0.783 0.701 0.749 0.891 0.721 0.855 0.737 0.757 0.831 0.726 0.755 0.722 0.737
336 0.903 0.730 0.748 0.849 0.639 0.766 0.691 0.714 0.810 0.693 0.695 0.749 0.715
5.2. Baselines
@@
In our comparison experiments, we adopted four popular baselines, including MLP, Transformer, and Mamba architectures. Informer (Zhou et al., 2021) introduces an efficient transformer for long sequence time-series forecasting using ProbSparse self-attention and self-attention distilling, enabling encoder-decoder architectures to handle long sequences effectively. MICN (Wang et al., 2023a) proposes a multi-scale context network that models both local and global contexts for long-term time series forecasting, capturing patterns across different temporal scales to enhance performance. FEDformer (Zhou et al., 2022) introduces a frequency-enhanced decomposed transformer that models time series in both time and frequency domains, improving long-term forecasting by effectively capturing temporal patterns. PatchTST (Nie et al., [n. d.]) applies transformers to time series by treating them as sequences of patches, enabling effective long-term forecasting through self-attention over patch representations to capture temporal dependencies.Autoformer (Wu et al., 2021), an earlier state-of-the-art model, leverages a decomposition architecture and auto-correlation mechanism to enhance efficiency and accuracy in long-term time series forecasting, outperforming traditional Transformer models. iTransformer (Liu et al., 2024a) is the latest and most effective Transformer-based model, utilizing attention and feed-forward networks on inverted dimensions, embedding time points into variate tokens. DLinear (Zeng et al., 2023) challenges the effectiveness of Transformer models by proposing a simple one-layer linear model that captures temporal relations in an ordered set of continuous points. It employs positional encoding and uses tokens to embed sub-series, preserving some ordering information in Transformers. Lastly, Mamba (Gu and Dao, 2023), a well-known sequential model from last year, uses a bidirectional Mamba block to extract inter-variate correlations and temporal dependencies. Additionally, we have selected five SOTA baselines  (Yu et al., 2018; Guo et al., 2019; Wu et al., 2019; Bai et al., 2020; Jiang et al., 2023) from traffic forecasting domain.
5.3. Implementation Details
@@
We adopted the default settings of the Time-Series-Library (Wu et al., 2022) to conduct a comprehensive comparison of baselines. We use the results of five random seeds as the average. We used 96 time steps as input and 336 time steps as ground truth. The code was implemented in PyTorch and executed on a V100 GPU with 32GB memory and 384GB RAM, provided by NCI Australia, an NCRIS-enabled capability supported by the Australian Government.
5.4. Results of Extremely Long Forecasting with gaps
@@
We use Mean Squared Error (MSE) and Mean Absolute Error (MAE) metrics to evaluate performance, averaging results across different seeds. It is observed that nearly all results are poor, highlighting the significant challenge posed by domain shifts over time for extremely long forecasting with gaps. These baseline results also indicate that traditional the-state-of-the-art (SOTA) rankings and methodologies are no longer effective. Notably, MICN, which performs the worst under conventional data and settings, shows the best performance in our setting. This is understandable because MICN’s core technology focuses on exploring the correlation within the data itself. This insight suggests that future efforts to tackle this problem should place greater emphasis on leveraging the intrinsic potential of the data. Additionally, we have compiled the training time for one epoch of all the baselines on the largest dataset, PEMS04_gap, and the smallest dataset, tfNSW, as shown in Table 9.
@@
Table 4. Results of ablation study with 4 different input step lengths
@@
5.5. Results of Hourly and Daily Forecasting
We believe that both the hourly and daily datasets are equally significant. Multi-scale, diverse datasets can provide the community with valuable references. We observe that the performance degrades progressively from the hourly to the daily to the gap datasets. Smaller time scales help reduce complexity and uncertainty, thereby improving the accuracy of prediction. Research has shown that clustering at different scales can improve model performance (Wang et al., 2024). Therefore, our aggregated version of the data will contribute new external features to the community.
Table 5. Comparison in hourly and daily datasets
@@ 5.5. Results of Hourly and Daily Forecasting
6. Prospects and Constraints
Prospects. Our dataset spans the longest time period among existing datasets, and it is not only the largest spatially, but also evolving in growth. It can continue to update according to the updates from the PeMS system in the future. It is specifically designed for various complex scenarios, such as those already mentioned, including extremely long forecasting with long gaps, and hourly and daily predictions.